OpenStudy (anonymous):

How do I solve: c^2=-15-8c

OpenStudy (anonymous):

x=-3,x=-5

OpenStudy (solomonzelman):

\(\Large\color{blue}{ c^2=-15-8c }\) add 8c to both sides, \(\Large\color{blue}{ c^2+8c=-15 }\) it is the same question as the last time, except that there the variable was z, and here it is c.

OpenStudy (salazarblack):

are you familiar with this formula? http://upload.wikimedia.org/math/5/6/4/5646d4fbfc6422d507dce0f4a17f74a8.png

OpenStudy (radar):

continuing: \[c^2 + 8c + 15 = 0\]

OpenStudy (radar):

Solve by factoring. It factors easily.

OpenStudy (solomonzelman):

or add 16 to both sides, when it was ...=-15 and complete the square .

OpenStudy (radar):

Completing the square also works nicely. Hey you have been provided three different ways to solve this. It is your choice.

OpenStudy (solomonzelman):

3??! quadratic formula is only a rescue way, when you can't factor, and when it is difficult to complete the square.. if you know what I mean....

OpenStudy (solomonzelman):

I would say that there are 2 ways, and another backup \(\Large\color{blue}{ :) }\)

OpenStudy (salazarblack):

but isn't it easier just to use the formula?

OpenStudy (salazarblack):

by factoring x1,x2 can be 3 and 5